OpenStudy (anonymous):

if log2(x)=log2(a)+log2(b)-log2(c) then what is x?

OpenStudy (anonymous):

\[\frac{ab}{c}\]

OpenStudy (nikita2):

x = (ab)/c. just use property of log: loga + logb = log(a*b) and log(a) - log(b) = log(a/b)